OpenStudy (anonymous):

I need to find 5 solutions to this problem...

OpenStudy (anonymous):

\[(x^2-5x+5)^{(x^2-9x+20)}=1 \] I attempted it so don't say "please try to attempt it" or something.

OpenStudy (anonymous):

I also know that any number with the power of zero is one.. but eh.

OpenStudy (tkhunny):

Are you SURE there are 5? Exponent is Zero but Base is NOT zero. \(x^{2} - 9x + 20 = 0\) \((x-5)(x-4) = 0\) \(x = 5\;or\;x = 4\) -- There's two. unless \(5^{2} - 5(5) + 5 = 0\;or\;4^{2} - 5(4) + 5 = 0\). Neither is zero, so we're good. Base is 1 but exponent is Finite \(x^{2} - 5x + 5 = 1\) \(x^{2} - 5x + 4 = 0\) \((x-4)(x-1)=0\) \(x = 4\;or\;x = 1\) -- Too bad. A repeat. We're up to 3!

OpenStudy (anonymous):

Yeah, there's 5 solutions unfortunately :(

OpenStudy (anonymous):

what happens if you try 2 ?

OpenStudy (anonymous):

if you get a base of \(-1\) and the exponent of something even, then it will work i think

OpenStudy (anonymous):

\[x^2-5x+5=-1\] \[x^2-5x+6=0\] \[(x-2)(x-3)=0\] so try \(x=2\) and see if the exponent is an even integer

OpenStudy (anonymous):

oh yeah it has be be even, because \(2\) is even so \(2\) will work

OpenStudy (tkhunny):

Both are strictly positive and greater than unity for x < 1 - There cannot be solutions out there. Both are strictly positive and greater than unity for x > \((1/2)(9+\sqrt{5}\) - There cannot be solutions out there. Anything interesting has to be in \([1,5.618]\)

OpenStudy (anonymous):

\(x=2\) works for sure base is \(-1\) an exponent is even

OpenStudy (anonymous):

maybe you will get lucky and 3 will work also base is \(-1\)

OpenStudy (anonymous):

yeah it works too

OpenStudy (anonymous):

5, 4, 1, 2 and 3? Those are the five solutions?

OpenStudy (tkhunny):

Very nice. It's not yet clear to me why I can't get my software to graph them. I pulled all the usual tricks for negative values and logarithms. It will take a little thought.

OpenStudy (anonymous):

Okay, I'm going to check them all right now.

OpenStudy (anonymous):

Wait.. I used 2 and got -1.. will that work?

OpenStudy (anonymous):

Thanks you guys, I wish that I could give both of you medals. :(